# Spatial exponential decay of perturbations in optimal control of general evolution equations

### v3

@@ 6.2. Domain-uniformly stabilizing control
$\begin{pmatrix}\xi_{1}\\ \xi_{2}\end{pmatrix}=\begin{pmatrix}\zeta_{1}+\zeta_{2}\\ \zeta_{1}-\zeta_{2}\end{pmatrix}=\begin{pmatrix}\frac{1}{c}&-\frac{1}{c}\\ 1&-1\end{pmatrix}\begin{pmatrix}\zeta_{1}\\ \zeta_{2}\end{pmatrix}=:T\zeta\implies\begin{pmatrix}\zeta_{1}\\ \zeta_{2}\end{pmatrix}=\begin{pmatrix}\frac{c}{2}&\frac{1}{2}\\ -\frac{c}{2}&\frac{1}{2}\end{pmatrix}\begin{pmatrix}\xi_{1}\\ \xi_{2}\end{pmatrix}=T^{-1}\xi,$
@@
$\begin{split}\forall\,(\omega,t)\in\Omega_{L}\times[0,T]:\frac{\partial}{\partial t}\zeta(\omega,t)&=\underset{=T^{-1}DT}{\underbrace{\begin{pmatrix}c&0\\ 0&-c\end{pmatrix}}}\frac{\partial}{\partial\omega}\zeta(\omega,t)-\underset{=T^{-1}FT}{\underbrace{\begin{pmatrix}k&0\\ 0&k\end{pmatrix}}}\chi_{\Omega_{L}^{c}}(\omega)\zeta(\omega,t)\\ \forall\,t\in[0,T]:\zeta_{1}(0,t)&=\zeta_{2}(0,t)\quad\textrm{and}\quad\zeta_{1}(L,t)=\zeta_{2}(L,t)\\ \forall\,\omega\in\Omega_{L}:\zeta_{1}(\omega,0)&=\frac{1}{2}\left(cx_{\Omega_{L}}^{0}(\omega)-\int_{0}^{\omega}x_{\Omega_{L}}^{1}(s)+k\,\chi_{\Omega_{L}^{c}}(s)x_{\Omega_{L}}^{0}(s)\mathrm{d}s\right)\\ \textrm{ and }\zeta_{2}(\omega,0)&=-\frac{1}{2}\left(cx_{\Omega_{L}}^{0}(\omega)+\int_{0}^{\omega}x_{\Omega_{L}}^{1}(s)+k\,\chi_{\Omega_{L}^{c}}(s)x_{\Omega_{L}}^{0}(s)\mathrm{d}s\right).\end{split}$ (54)
@@
$\zeta_{1}(\omega,0)=\left\{\begin{array}[]{cc}\frac{1}{2}\left(cx_{\Omega_{L}}^{0}(\omega)-\int_{0}^{\omega}x_{\Omega_{L}}^{1}(s)+k\,\chi_{\Omega_{L}^{c}}(s)x_{\Omega_{L}}^{0}(s)\mathrm{d}s\right),&\omega\in[0,L]\\ -\frac{1}{2}\left(cx_{\Omega_{L}}^{0}(2L-\omega)+\int_{0}^{2L-\omega}x_{\Omega_{L}}^{1}(s)+k\,\chi_{\Omega_{L}^{c}}(s)x_{\Omega_{L}}^{0}(s)\mathrm{d}s\right),&\omega\in[L,2L]\end{array}\right$.
@@ Appendix B Proof of Lemmata 21-23
This shows (23). Checking the initial and boundary conditions is straightforward. The solution formula can be extended to $x^{0}\in L^{2}([0,L])$ via a standard density argument (see [25, Proposition II.1.5.]). ∎
@@ Appendix C Proof of Lemma 27
Now Gronwall’s Lemma [15, Lemma 2.4] yields $p_{1}(t)=p_{2}(t)$ for all $t\in(0,t_{2})$ and therefore a contradiction. An analogous argumentation yields the result for (38). ∎
@@
Again the solution formula can be extended to $x^{0}\in L^{2}([0,L])$ via (see [25, Proposition II.1.5.]). ∎
@@ Appendix D Proof of Proposition 31
$\frac{\partial}{\partial t}x(t)-\frac{\partial}{\partial t}x(0)+2k\,\chi_{\Omega_{L}^{c}}(x(t)-x(0))+\int_{0}^{t}k^{2}\,\chi_{\Omega_{L}^{c}}x(\tau)\mathrm{d}\tau=-\frac{\partial}{\partial\omega}v(t)-\frac{\partial}{\partial t}x(0)-k\,\chi_{\Omega_{L}^{c}}x(0)-k\,\chi_{\Omega_{L}^{c}}\int_{0}^{t}\frac{\partial}{\partial\omega}v(\tau)\mathrm{d}\tau.$

### v4

@@ 6.2. Domain-uniformly stabilizing control
$\begin{pmatrix}\xi_{1}\\ \xi_{2}\end{pmatrix}=\begin{pmatrix}\frac{1}{c}\left(\zeta_{1}+\zeta_{2}\right)\\ \zeta_{1}-\zeta_{2}\end{pmatrix}=\begin{pmatrix}\frac{1}{c}&\frac{1}{c}\\ 1&-1\end{pmatrix}\begin{pmatrix}\zeta_{1}\\ \zeta_{2}\end{pmatrix}=:T\zeta\implies\begin{pmatrix}\zeta_{1}\\ \zeta_{2}\end{pmatrix}=\begin{pmatrix}\frac{c}{2}&\frac{1}{2}\\ \frac{c}{2}&-\frac{1}{2}\end{pmatrix}\begin{pmatrix}\xi_{1}\\ \xi_{2}\end{pmatrix}=T^{-1}\xi,$
@@
$\begin{split}\forall\,(\omega,t)\in\Omega_{L}\times[0,T]:\frac{\partial}{\partial t}\zeta(\omega,t)&=\underset{=T^{-1}DT}{\underbrace{\begin{pmatrix}-c&0\\ 0&c\end{pmatrix}}}\frac{\partial}{\partial\omega}\zeta(\omega,t)-\underset{=T^{-1}FT}{\underbrace{\begin{pmatrix}k&0\\ 0&k\end{pmatrix}}}\chi_{\Omega_{L}^{c}}(\omega)\zeta(\omega,t)\\ \forall\,t\in[0,T]:\zeta_{1}(0,t)&=\zeta_{2}(0,t)\quad\textrm{and}\quad\zeta_{1}(L,t)=\zeta_{2}(L,t)\\ \forall\,\omega\in\Omega_{L}:\zeta_{1}(\omega,0)&=\frac{1}{2}\left(cx_{\Omega_{L}}^{0}(\omega)-\int_{0}^{\omega}x_{\Omega_{L}}^{1}(s)+k\,\chi_{\Omega_{L}^{c}}(s)x_{\Omega_{L}}^{0}(s)\mathrm{d}s\right)\\ \textrm{ and }\zeta_{2}(\omega,0)&=\frac{1}{2}\left(cx_{\Omega_{L}}^{0}(\omega)+\int_{0}^{\omega}x_{\Omega_{L}}^{1}(s)+k\,\chi_{\Omega_{L}^{c}}(s)x_{\Omega_{L}}^{0}(s)\mathrm{d}s\right).\end{split}$ (54)
@@
$\zeta_{1}(\omega,0)=\left\{\begin{array}[]{cc}\frac{1}{2}\left(cx_{\Omega_{L}}^{0}(\omega)-\int_{0}^{\omega}x_{\Omega_{L}}^{1}(s)+k\,\chi_{\Omega_{L}^{c}}(s)x_{\Omega_{L}}^{0}(s)\mathrm{d}s\right),&\omega\in[0,L]\\ \frac{1}{2}\left(cx_{\Omega_{L}}^{0}(2L-\omega)+\int_{0}^{2L-\omega}x_{\Omega_{L}}^{1}(s)+k\,\chi_{\Omega_{L}^{c}}(s)x_{\Omega_{L}}^{0}(s)\mathrm{d}s\right),&\omega\in[L,2L]\end{array}\right$.
@@ Appendix B Proof of Lemmata 21-23
This shows (23). Checking the initial and boundary conditions is straightforward. The solution formula can be extended to $x^{0}\in L^{2}([0,L])$ via a standard density argument (see [23, Proposition II.1.5.]). ∎
@@ Appendix C Proof of Lemma 27
Now Gronwall’s Lemma [14, Lemma 2.4] yields $p_{1}(t)=p_{2}(t)$ for all $t\in(0,t_{2})$ and therefore a contradiction. An analogous argumentation yields the result for (38). ∎
@@
Again the solution formula can be extended to $x^{0}\in L^{2}([0,L])$ via (see [23, Proposition II.1.5.]). ∎
@@ Appendix D Proof of Proposition 31
$\frac{\partial}{\partial t}x(t)-\frac{\partial}{\partial t}x(0)+2k\,\chi_{\Omega_{L}^{c}}(x(t)-x(0))+\int_{0}^{t}k^{2}\,\chi_{\Omega_{L}^{c}}x(\tau)\mathrm{d}\tau$
$=$ $-\frac{\partial}{\partial\omega}v(t)-\frac{\partial}{\partial t}x(0)-k\,\chi_{\Omega_{L}^{c}}x(0)-k\,\chi_{\Omega_{L}^{c}}\int_{0}^{t}\frac{\partial}{\partial\omega}v(\tau)\mathrm{d}\tau.$